OpenStudy (anonymous):

What values of x and y satisfy x − 3y = 4 and 2x + y = 7 simultaneously? A) \[x=\frac{ 1 }{ 7 }~and~x=\frac{ 25 }{ 7 }\] B) \[x=\frac{ 25 }{ 7 } ~and~ y=−\frac{ 1 }{ 7 }\] C) x=7 and y=1 D) x=3 and y=1 @kelliegirl33 @agent0smith @AllTehMaffs @jdoe0001 @RICARDOismyfirstname @shrutipande9

OpenStudy (anonymous):

these are a pain in the butt.....we are going to have to sub in your answer choices to see which one satisfies both equations...

OpenStudy (anonymous):

lets just solve it and see if that is an answer choice....hold on

OpenStudy (anonymous):

If you just want the answer then its option 'b'

OpenStudy (anonymous):

x - 3y = 4 -->(-2) 2x + y = 7 -------------- -2x + 6y = -8 2x + y = 7 ------------- 7y = -1 y = -1/7 x - 3y = 4 x - 3(-1/7) = 4 x + 3/7 = 4 x = 4 - 3/7 x = 28/7 - 3/7 x = 25/7 found it.....x = 25/7 and y = -1/7

OpenStudy (anonymous):

Thanks :)